OpenStudy (anonymous):

What is the exact value of the expression sqrt(9y^4) + 12*sqrt(y^6) - 3*sqrt(y^6) - sqrt(y^4)

OpenStudy (anonymous):

Please help!! Medals for everyone that does!

OpenStudy (anonymous):

okay first you have to simplify the square roots and then combine like terms. so it goes like this: \[\sqrt{9y^4}+12\sqrt{y^6}-3\sqrt{y^6}-\sqrt{y^4}\] = \[3y^2+12y^3-3y^3-y^2\] = \[9y^3+2y^2\]

OpenStudy (anonymous):

That wasn't any of the possible answers: 8y^2 + 9y^3 11y^2 11y^5 2y^2 + 9y^3

OpenStudy (anonymous):

@GuardianOfLore

OpenStudy (anonymous):

the last option is the same, you're just switching the terms around. its called the commutative property. the algebraic rule is that A+B=B+A. so, if i have 2y^2 + 9y^3, i can switch the terms around and have 9y^3 + 2y^2. is that helpful?

OpenStudy (anonymous):

Yes it is, thank you, I didn't look hard enough, thank you for all of your help!

OpenStudy (anonymous):

its all good man :) anytime.